# A Deep Learning Method for Simultaneous Denoising and Missing Wedge Reconstruction in Cryogenic Electron Tomography

### v2

@@ Acknowledgements
Appendix A Background on Self-Supervised Deep Learning Techniques
The DeepDeWedge loss is inspired by noise2noise and noisier2noise self-supervised learning. Here, we give a brief overview of the main ideas behind these two frameworks.
A.1 Denoising with Noise2Noise
@@
Noise2noise-based methods also aim to train a neural network to map a noisy image to a clean one. Contrary to supervised learning, noise2noise assumes access to a dataset of pairs of noisy observations ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{y}^{0}_{i}}=\mathbf{x}^{*}_{i}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{0}_{i}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{y}^{1}_{i}}=\mathbf{x}^{*}_{i}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{n}^{1}_{i}}$ for each ground-truth image $\mathbf{x}^{*}_{i}$. Moreover, it assumes that the noise terms ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{n}^{0}_{i}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{n}^{1}_{i}}$ are independent and that the noise ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{n}^{1}_{i}}$ is zero-mean. Then, one can train a network $f_{\text{$\boldsymbol{\theta}$}}$ for denoising to map the noisy observation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{y}^{0}}$ onto its counterpart ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{y}^{1}}$. Formally, one can show that
@@
$\mathop{\mathbb{E}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{y}^{0}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{y}^{1}}}\Big{[}\left|\left|f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{y}^{0}})-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{y}^{1}}\right|\right|_{2}^{2}\Big{]}=\mathop{\mathbb{E}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{y}^{0}},\mathbf{x}^{*}}\Big{[}\left|\left|f_{\text{$\boldsymbol{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{y}^{0}})-\mathbf{x}^{*}\right|\right|_{2}^{2}\Big{]}+c,$ (8)
@@
where $c$ is a constant that is independent of the network weights $\boldsymbol{\theta}$. Thus, a self-supervised noise2noise training objective based on samples approximates, up to an additive constant, the same underlying risk (left side of Equation (8)) as the supervised objective function. This approximation becomes better as the number of training examples increases and theoretically and empirically, networks trained with a noise2noise-like loss perform as well as if trained on sufficiently many examples [KAH23].
A.2 Denoising and Recovering Missing Data with Noiser2Noise
@@
IsoNet’s denoising approach is motivated by the noisier2noise framework [Mor+20]. While training a model for image denoising with a noise2noise objective requires paired noisy observations, a noisier2noise objective is based on a single noisy observation per image.
@@
Noisier2Noise for Denoising: We assume that we have access to a dataset of noisy images. Contrary to noise2noise, we assume that we have only one single noisy observation $\mathbf{y}=\mathbf{x}^{*}+\mathbf{n}$ per ground-truth $\mathbf{x}^{*}$. The goal is to train a neural network for image denoising. For constructing a training objective with noisier2noise, we assume that the noise terms $\mathbf{n}$ that corrupt all images come from the same distribution and that we are able to sample from this distribution. We construct model inputs $\tilde{\mathbf{y}}$ by further corrupting the noisy images $\mathbf{y}$ with an additional noise term $\tilde{\mathbf{n}}$ sampled from the true noise distribution, i.e., $\tilde{\mathbf{y}}=\mathbf{y}+\tilde{\mathbf{n}}$. [Mor+20, ] showed that if we train a neural network $f_{\text{$\boldsymbol{\theta}$}}$ to map these noisier model inputs $\tilde{\mathbf{y}}$ onto their single noisy counterparts $\mathbf{y}$ using the squared L2-loss, one can use the final trained network to construct a denoiser for the double noisy images $\tilde{\mathbf{y}}$.
@@
Noisier2Noise for Missing Data Recovery: We use ideas from noisier2noise for missing wedge reconstruction. To illustrate how noisier2noise can be used to predict missing data, consider the following setup: Assume we have a dataset of masked measurements $\mathbf{y}=\mathbf{M}\mathbf{x}^{*}$. Here, $\mathbf{M}$ denotes a random mask, i.e., a diagonal matrix with entries in $\{0,1\}$, and $\mathbf{x}^{*}$ is an unknown ground-truth signal vector. Like above, the ground truths and masks are different for each measurement $\mathbf{y}$ and the dataset contains only the measurements and no ground truths. In this setup, we want to train a neural network $f_{\text{$\boldsymbol{\theta}$}}$ to predict the data $\mathbf{x}^{*}$ from the measurement $\mathbf{y}$. To do this, we construct double-masked measurements $\tilde{\mathbf{y}}=\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{y}$, where $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ is another random mask, and train the network to map the double-masked measurements $\tilde{\mathbf{y}}$ to their counterparts $\mathbf{y}$. It can be shown that, under certain assumptions on the random masks $\mathbf{M}$ and $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$, the final trained model can be used in an algorithm to recover the data missing in the singly masked measurements $\mathbf{y}$. As for denoising, the model estimates the missing data based on the doubly-masked measurement $\tilde{\mathbf{y}}$. For details, we refer readers to recent works by [MC23][MC23, MC22], who discussed noisier2noise for the recovery of missing data in accelerated magnetic resonance imaging.
Appendix B Normalizing Tomograms and Sub-Tomograms
@@
During model fitting, we apply Gaussian normalization to each model input sub-tomogram ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}$ by subtracting a global mean $\mu$ and dividing by a global standard deviation $\sigma$, i.e.,
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\leftarrow\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}-\mu}{\sigma}.$ (9)
@@
$\mu=\frac{1}{M}\sum_{i=1}^{M}\text{mean}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\right),\qquad\sigma^{2}=\frac{1}{M}\sum_{i=1}^{M}\text{var}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\boldsymbol{\varphi}$}_{i}}}^{\text{0}}}\right).$ (10)
@@
Assume we want to obtain the DeepDeWedge reconstruction of a tilt series $\mathbf{t}$. For this, we have to apply a fitted model to sub-tomograms of the FBP reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})$, where ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}}$ are the result of splitting the tilt series $\mathbf{t}$ using the even/odd or frame-based split. Before extracting sub-tomograms from the reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})$, we normalize them to have a mean $\mu_{\mathbf{t}}$ and variance $\sigma^{2}_{\mathbf{t}}$ via
@@
$\overline{\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})}$ $=\frac{\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})-\text{mean}\big{(}\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})\big{)}}{\sqrt{\text{var}\big{(}\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{t}^{\text{0}}})\big{)}}}\cdot\sigma_{\mathbf{t}}+\mu_{\mathbf{t}},$ (11)
@@
$\overline{\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})}$ $=\frac{\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})-\text{mean}\big{(}\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})\big{)}}{\sqrt{\text{var}\big{(}\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathbf{t}^{\text{1}}})\big{)}}}\cdot\sigma_{\mathbf{t}}+\mu_{\mathbf{t}}.$ (12)

### v3

@@ Acknowledgements
Supplementary Information
Appendix 1 Background on Self-Supervised Deep Learning Techniques
The DeepDeWedge loss is inspired by Noise2Noise [10] and Noisier2Noise [16] self-supervised learning. Here, we give a brief overview of the main ideas behind these two frameworks.
1.1 Denoising with Noise2Noise
@@
Noise2Noise-based methods also aim to train a neural network to map a noisy image to a clean one. Contrary to supervised learning, Noise2Noise assumes access to a dataset of pairs of noisy observations ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{y}^{0}_{i}}=\mathbf{x}^{*}_{i}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{0}_{i}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}^{1}_{i}}=\mathbf{x}^{*}_{i}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{n}^{1}_{i}}$ for each ground-truth image $\mathbf{x}^{*}_{i}$. Moreover, it assumes that the noise terms ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{n}^{0}_{i}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{n}^{1}_{i}}$ are independent and that the noise ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{n}^{1}_{i}}$ is zero-mean. Then, one can train a network $\mathrm{f}_{\text{$\bm{\theta}$}}$ for denoising to map the noisy observation ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{y}^{0}}$ onto its counterpart ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}^{1}}$. Formally, one can show that
@@
$\mathop{\mathbb{E}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{y}^{0}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}^{1}}}\Big{[}\left|\left|\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{y}^{0}})-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}^{1}}\right|\right|_{2}^{2}\Big{]}=\mathop{\mathbb{E}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{y}^{0}},\mathbf{x}^{*}}\Big{[}\left|\left|\mathrm{f}_{\text{$\bm{\theta}$}}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{y}^{0}})-\mathbf{x}^{*}\right|\right|_{2}^{2}\Big{]}+c,$ (8)
@@
where $c$ is a constant that is independent of the network weights $\bm{\theta}$. Thus, a self-supervised Noise2Noise training objective based on samples approximates, up to an additive constant, the same underlying risk (left side of Equation (8)) as the supervised objective function. This approximation becomes better as the number of training examples increases and theoretically and empirically, networks trained with a Noise2Noise-like loss perform as well as if trained on sufficiently many examples [33].
1.2 Denoising and Recovering Missing Data with Noiser2Noise
@@
IsoNet’s denoising approach is motivated by the Noisier2Noise framework [16]. While training a model for image denoising with a Noise2Noise objective requires paired noisy observations, a Noisier2Noise objective is based on a single noisy observation per image.
@@
Noisier2Noise for Denoising: We assume that we have access to a dataset of noisy images. Contrary to Noise2Noise, we assume that we have only one single noisy observation $\mathbf{y}=\mathbf{x}^{*}+\mathbf{n}$ per ground-truth $\mathbf{x}^{*}$. The goal is to train a neural network for image denoising. For constructing a training objective with Noisier2Noise, we assume that the noise terms $\mathbf{n}$ that corrupt all images come from the same distribution and that we are able to sample from this distribution. We construct model inputs $\tilde{\mathbf{y}}$ by further corrupting the noisy images $\mathbf{y}$ with an additional noise term $\tilde{\mathbf{n}}$ sampled from the true noise distribution, i.e., $\tilde{\mathbf{y}}=\mathbf{y}+\tilde{\mathbf{n}}$. Moran et al. [16] showed that if we train a neural network $\mathrm{f}_{\text{$\bm{\theta}$}}$ to map these noisier model inputs $\tilde{\mathbf{y}}$ onto their single noisy counterparts $\mathbf{y}$ using the squared L2-loss, one can use the final trained network to construct a denoiser for the double noisy images $\tilde{\mathbf{y}}$.
@@
Noisier2Noise for Missing Data Recovery: We use ideas from Noisier2Noise for missing wedge reconstruction. To illustrate how Noisier2Noise can be used to predict missing data, consider the following setup: Assume we have a dataset of masked measurements $\mathbf{y}=\mathbf{M}\mathbf{x}^{*}$. Here, $\mathbf{M}$ denotes a random mask, i.e., a diagonal matrix with entries in $\{0,1\}$, and $\mathbf{x}^{*}$ is an unknown ground-truth signal vector. Like above, the ground truths and masks are different for each measurement $\mathbf{y}$ and the dataset contains only the measurements and no ground truths. In this setup, we want to train a neural network $\mathrm{f}_{\text{$\bm{\theta}$}}$ to predict the data $\mathbf{x}^{*}$ from the measurement $\mathbf{y}$. To do this, we construct double-masked measurements $\tilde{\mathbf{y}}=\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}\mathbf{y}$, where $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$ is another random mask, and train the network to map the double-masked measurements $\tilde{\mathbf{y}}$ to their counterparts $\mathbf{y}$. It can be shown that, under certain assumptions on the random masks $\mathbf{M}$ and $\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{M}}}$, the final trained model can be used in an algorithm to recover the data missing in the singly masked measurements $\mathbf{y}$. As for denoising, the model estimates the missing data based on the doubly-masked measurement $\tilde{\mathbf{y}}$. For details, we refer readers to recent works by Millard and Chiew [34, 35], who discussed Noisier2Noise for the recovery of missing data in accelerated magnetic resonance imaging.
Appendix 2 Normalizing Tomograms and Sub-Tomograms
@@
During model fitting, we apply Gaussian normalization to each model input sub-tomogram ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}$ by subtracting a global mean $\mu$ and dividing by a global standard deviation $\sigma$, i.e.,
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\leftarrow\frac{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}-\mu}{\sigma}.$ (9)
@@
$\mu=\frac{1}{M}\sum_{i=1}^{M}\text{mean}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\right),\qquad\sigma^{2}=\frac{1}{M}\sum_{i=1}^{M}\text{var}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\mathbf{v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i,\text{$\bm{\varphi}$}_{i}}}^{\text{0}}}\right).$ (10)
@@
Assume we want to obtain the DeepDeWedge reconstruction of a tilt series $\mathbf{t}$. For this, we have to apply a fitted model to sub-tomograms of the FBP reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})$, where ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}}$ and ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}}$ are the result of splitting the tilt series $\mathbf{t}$ using the even/odd or frame-based split. Before extracting sub-tomograms from the reconstructions $\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})$ and $\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})$, we normalize them to have a mean $\mu_{\mathbf{t}}$ and variance $\sigma^{2}_{\mathbf{t}}$ via
@@
$\overline{\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})}$ $=\frac{\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})-\text{mean}\big{(}\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})\big{)}}{\sqrt{\text{var}\big{(}\text{FBP}({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{t}^{\text{0}}})\big{)}}}\cdot\sigma_{\mathbf{t}}+\mu_{\mathbf{t}},$ (11)
@@
$\overline{\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})}$ $=\frac{\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})-\text{mean}\big{(}\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})\big{)}}{\sqrt{\text{var}\big{(}\text{FBP}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{t}^{\text{1}}})\big{)}}}\cdot\sigma_{\mathbf{t}}+\mu_{\mathbf{t}}.$ (12)